# Rotatable IRS Aided Wireless Communication

### v1

@@ I Introduction
The development of the sixth-generation (6G) wireless networks is driven by the ambition to achieve unprecedented data rates, superior energy efficiency, and broader network coverage, which aim to support the demanding requirements of future ubiquitous connectivity [wu_6G, AIRS_magazine, xue_survey]. As a promising solution, intelligent reflecting surfaces (IRSs) have been introduced as a cost- and energy-efficient complement to traditional wireless networks, offering proactive and flexible control of the radio-wave propagation. By dynamically adjusting the amplitude and/or phase of passive elements, IRSs can reconfigure wireless channels to strengthen desired signals, suppress interference, improve physical-layer security, and extend network coverage in challenging propagation scenarios. These capabilities allow the IRS to support diverse applications such as interference mitigation in dense networks, throughput improvement for cell-edge users, and simultaneous wireless information and power transfer for energy-limited devices [CI, ISI, chen_iot]. Motivated by these prospects, extensive research has been conducted on IRS-aided systems, focusing on main issues including energy efficiency maximization, joint active and passive beamforming design, and the theoretical analysis of achievable rate regions [hirs_peng, kang_deployment, fixed_feng, fu_deployment].
@@
The aforementioned studies mainly consider IRSs that are installed at fixed locations with pre-determined orientations to enhance wireless communication systems. In existing works, the IRS is primarily deployed in terrestrial scenarios, where it is typically mounted on the outer walls of buildings, indoor ceilings, and other static structures [double_IRS, peng_deployment, magazine]. As a fixed IRS can only provide effective service to transmitters and receivers located within the same half-space relative to its reflective surface, communication nodes positioned on the opposite sides of the IRS cannot benefit from its reflection capability, which restricts its spatial coverage. Moreover, although passive beamforming can be used to reflect signals toward users, each reflection may lead to partial scattering of the electromagnetic waves in undesired directions. This dispersion of signal energy results in additional propagation loss and severe attenuation, which significantly degrade the overall signal quality and reduce the efficiency of the IRS-aided links, especially in complex propagation environments.
@@
To address the limitations of conventional fixed IRSs, recent research has explored dynamic deployment strategies that can operate in a more adaptive and flexible manner. These strategies introduce movable and rotatable IRS in wireless communication systems, thereby opening new opportunities for fully harnessing the potential of the IRS technology in next-generation wireless systems. To be specific, aerial IRS–enabled three-dimensional (3D) wireless networks, where the IRS is mounted on aerial platforms, e.g., unmanned aerial vehicles (UAVs), have received significant attention for intelligent signal reflection [lu_distance, liu_uav, uav_jiang, uav_wang]. Different from traditional terrestrial IRSs, which are inherently restricted to serving only the nodes located within the same half-space of the reflective surface, an aerial IRS can achieve near-omnidirectional reflection and facilitate line-of-sight (LoS) communication between transmitters and receivers due to its elevated position relative to the ground. In [lu_distance], transmit beamforming, IRS placement, and 3D passive beamforming were jointly optimized to maximize the worst-case signal-to-noise ratio (SNR) within a target area, utilizing a beam broadening and flattening technique to achieve significant performance gains over the fixed IRS. Furthermore, the maneuverability of UAVs allows the location and orientation of an aerial IRS to be dynamically adjusted, which introduces additional degrees of freedom for system optimization and enables highly flexible 3D network design that can adapt to real-time changes in communication demands. The work [liu_uav] developed an efficient alternating optimization algorithm in a UAV-mounted IRS-aided multicast system with the joint design of the IRS’s 3D location, 3D orientation, and passive beamforming. Despite many advantages mentioned above, aerial IRS poses fundamental limitations for several reasons. First, the endurance, stability, and controllability of the aerial platform remain critical technical concerns, as the platform must support sustained operation under varying environmental conditions [endurance]. Second, platform drift and vibration, which influence reflection accuracy and system stability, as well as the complexity associated with optimizing the placement and trajectory of aerial platforms, impose significant design burdens [complexity]. Third, regulatory constraints related to flight zones and airspace management may further limit the large-scale or long-term deployment of aerial IRSs in practical scenarios [no-fly].
@@
Considering the ease of implementation provided by the fixed IRS and the high flexibility offered by the aerial IRS, the rotatable IRS has been proposed, which focuses on dynamic orientation adjustment rather than full spatial displacement [rotate4, rotate1, rotate2, rotate3, rotate7, 6DMA1]. To be specific, a rotatable IRS can flexibly vary its rotation angles at a fixed position, thereby dynamically reconfiguring its reception and reflection patterns and providing additional spatial degrees of freedom without physical movement. By properly adjusting its rotation angle, a single rotatable IRS can achieve panoramic coverage or full-angle reflection, enabling it to serve any transmitter-receiver pair within its coverage region. Moreover, the practical implementation of a rotatable IRS resembles existing rotatable antennas, where the IRS is mounted on a motor-driven shaft, and its rotation angles is precisely controlled through mechanical adjustment mechanisms, which can also be considered as a special case of the general six-dimensional movable antenna (6DMA) model with both 3D position and 3D orientation control [6DMA2, 6DMA3, 6DMA4, RA_zheng, FR_lu, ning]. Such an architecture ensures strong compatibility with existing wireless infrastructures while avoiding the need for continuous positioning, complex mechanical control, and the high power consumption typically associated with mobile platforms. Because of its simple structure, low energy cost, and low latency, a rotatable IRS is appealing for scenarios that require rapid adaptation but where large-scale mechanical movement is impractical. The rotatable IRS has been extensively investigated under various communication systems, including cellular systems [rotate1], mobile edge computing systems [rotate2], and multicast systems [rotate3]. In [rotate1], the received power was maximized when the incident angle was zero. The work [rotate2] proposed a deep learning-based algorithm for the joint design of IRS orientation, phase shifts, and offloading strategies. The base station (BS) beamforming, IRS phase shifts, and orientation adjustment were jointly optimized to maximize the multicast rates in [rotate2].
@@
Despite the effectiveness of the IRS rotation strategies, most existing studies have relied on a simplified and idealized channel model that primarily accounts for the propagation path loss and orientation parameters, while overlooking other crucial physical-layer characteristics such as the effective reception area of the IRS, as well as its reception and reflection efficiency [physical1, physical2, physical3]. Consequently, the corresponding deployment strategies may underperform in practical scenarios with non-ideal channels, as they fail to capture the geometric and physical constraints inherent in practical channel conditions [PL]. To address this issue, the work [rotate5] assumed an IRS with predetermined electromagnetic properties and focused on how its steering capability affects the spatial power redistribution of the reflected beam. Furthermore, a physics‑based model was proposed in [rotate6], where the IRS elements were grouped into tiles whose effects on the wireless channel were characterized by solving the electric and magnetic field integral equations. In the above studies, the IRS is represented as a continuous reflecting surface whose global response determines the scattered beam. It provides analytical simplicity and facilitates theoretical investigations. In contrast, the IRS can be modeled as a planar distribution of discrete controllable elements that capture the collective response produced by the individual reflecting elements, which provides a more accurate and practical framework for IRS design by capturing its physical characteristics. One of the main advantages lies in allowing the electromagnetic properties to be directly integrated with circuit‑level models of the tunable elements, thereby bridging the gap between idealized analysis and practical design. In [cos1], the authors evaluated the scattering performance, power, and size of a single IRS element and then maximized the received signal power toward a specific user. Furthermore, the rotation of the IRS was optimized to maximize the worst SNR over all locations in a target area in both indoor [cos2] and outdoor scenarios [cos3]. However, in [cos1, cos2, cos3], the radiation patterns of the transmit and receive antennas, as well as the IRS elements, are modeled with angle‑dependent factors derived from a simplified cosine law, which may lead to limitations of rotation optimization in terms of SNR.
@@ II System Model
As shown in Fig. 1, we consider an IRS-aided wireless communication system, where the IRS is deployed to assist from the multi-antenna BS to the single-antenna user within the target area $\mathcal{A}$. We focus on a challenging scenario where the direct BS-user link is blocked by obstacles and assume that all the considered channels follow the free-space LoS model [LOS1, LOS2]. The IRS consists of a total of $N=N_{\mathrm{c}}\times N_{\mathrm{r}}$ elements, arranged in $N_{\mathrm{r}}$ rows and $N_{\mathrm{c}}$ columns, with each element have the size of $\bar{l}\times\bar{l}$. The BS is equipped with a conventional UPA placed on the $x$-$y$ plane, and the total number of antennas is $M=M_{\mathrm{c}}\times M_{\mathrm{r}}$ antennas, where $M_{\mathrm{r}}$ and $M_{\mathrm{c}}$ denote the number of antennas along the $x$- and $y$-axis, respectively. In the considered system, we establish a global Cartesian coordinate system (CCS). Assuming that the target area $\mathcal{A}$ is a rectangular area with length $A_{\mathrm{x}}$ and width $A_{\mathrm{y}}$ on the $x$-$y$ plane, its center is denoted as $\bm{p}_{\mathrm{r}}=\left[x_{\mathrm{r}},y_{\mathrm{r}},0\right]^{T}$. Thus, any user location within the target area $\mathcal{A}$ is located at $\bm{p}_{\mathrm{U}}=\left[x_{\mathrm{U}},y_{\mathrm{U}},0\right]^{T}$ with $x_{\mathrm{U}}\in\left[x_{\mathrm{r}}-A_{\mathrm{x}}/2,x_{\mathrm{r}}+A_{\mathrm{x}}/2\right]$ and $y_{\mathrm{U}}\in\left[y_{\mathrm{r}}-A_{\mathrm{y}}/2,y_{\mathrm{r}}+A_{\mathrm{y}}/2\right]$.
@@
We denote $\varpi_{\mathrm{B}}^{\mathrm{c}}=\arccos\left({\frac{{({{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}})\cdot{\bm{m}}_{\mathrm{B}}^{\mathrm{c}}}}{{\|{{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}}\|\cdot\|{{\bm{m}}_{\mathrm{B}}^{\mathrm{c}}}\|}}}\right)$, $\varpi_{\mathrm{B}}^{\mathrm{r}}=\arccos\left({\frac{{({{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}})\cdot{\bm{m}}_{\mathrm{B}}^{\mathrm{r}}}}{{\|{{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}}\|\cdot\|{{\bm{m}}_{\mathrm{B}}^{\mathrm{r}}}\|}}}\right)$, ${\varpi_{\mathrm{c}}}=\arccos\left({\frac{{({{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}})\cdot{\bm{m}}_{\mathrm{c}}}}{{\|{{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}}\|\cdot\|{{\bm{m}}_{\mathrm{c}}}\|}}}\right)$, and ${\varpi_{\mathrm{r}}}=\arccos\left({\frac{{({{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}})\cdot{\bm{m}}_{\mathrm{r}}}}{{\|{{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}}\|\cdot\|{{\bm{m}}_{\mathrm{r}}}\|}}}\right)$ as the angles between $({{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}})$ and ${\bm{m}}_{\mathrm{B}}^{\mathrm{c}}$}, ${\bm{m}}_{\mathrm{B}}^{\mathrm{r}}$}, ${\bm{m}}_{\mathrm{c}}$}}, and ${\bm{m}}_{\mathrm{r}}$}}, respectively. According to [double_IRS], the LoS channel from the BS to the IRS can be assumed to be a rank-one matrix if $t_{1,1}\geq\sqrt{N}l^{2}/\lambda$. To facilitate the design of the BS and IRS beamforming, we provide a tractable characterization for the BS-IRS channel in the following proposition.
@@
Given $\bm{\Phi}$ and $\bm{\Omega}$, the optimal transmit beamforming at the BS is given by [lu_distance]
@@ VI Conclusion
We consider a perfectly conducting rectangular plate with dimensions $l_{1}\times l_{2}$ and negligible thickness, located in the xyx-y plane. A distant point source radiates linearly polarized waves with wave number $k=\frac{2\pi}{\lambda}$. As shown in Fig. 13, let $\theta^{\mathrm{i}}$, $\phi^{\mathrm{i}}$, $\theta^{\mathrm{r}}$, and $\phi^{\mathrm{r}}$ denote the azimuth and elevation angles of the incident wave of the center of plane, as well as the elevation and azimuth angles of the reflected wave of the center of plane. The electric and magnetic field components of the incident wave is given by [EM1, EM2]

### v2

# arXiv:2511.10006v2[eess.SP] 03 Mar 2026

@@ I Introduction
The development of the sixth-generation (6G) wireless networks is driven by the ambition to achieve unprecedented data rates, superior energy efficiency, and broader network coverage, which aim to support the demanding requirements of future ubiquitous connectivity [42, 17, 45]. As a promising solution, intelligent reflecting surfaces (IRSs) have been introduced as a cost- and energy-efficient complement to traditional wireless networks, offering proactive and flexible control of the radio-wave propagation [47, 12, 11]. By dynamically adjusting the amplitude and/or phase of passive elements, IRSs can reconfigure wireless channels to strengthen desired signals, suppress interference, improve physical-layer security, and extend network coverage in challenging propagation scenarios. These capabilities allow the IRS to support diverse applications such as interference mitigation in dense networks, throughput improvement for cell-edge users, and simultaneous wireless information and power transfer for energy-limited devices [29, 35, 1]. Motivated by these prospects, extensive research has been conducted on IRS-aided systems, focusing on main issues including energy efficiency maximization, joint active and passive beamforming design, and the theoretical analysis of achievable rate regions [30, 16, 5, 6].
@@
The aforementioned studies mainly consider IRSs that are installed at fixed locations with pre-determined orientations to enhance wireless communication systems. In existing works, the IRS is primarily deployed in terrestrial scenarios, where it is typically mounted on the outer walls of buildings, indoor ceilings, and other static structures [9, 31, 43]. As a fixed IRS can only provide effective service to transmitters and receivers located within the same half-space relative to its reflective surface, communication nodes positioned on the opposite sides of the IRS cannot benefit from its reflection capability, which restricts its spatial coverage. Moreover, although passive beamforming can be used to reflect signals toward users, each reflection may lead to partial scattering of the electromagnetic waves in undesired directions. This dispersion of signal energy results in additional propagation loss and severe attenuation, which significantly degrade the overall signal quality and reduce the efficiency of the IRS-aided links, especially in complex propagation environments.
@@
To address the limitations of conventional fixed IRSs, recent research has explored dynamic deployment strategies that can operate in a more adaptive and flexible manner. These strategies introduce movable and rotatable IRS in wireless communication systems, thereby opening new opportunities for fully harnessing the potential of the IRS technology in next-generation wireless systems. To be specific, aerial IRS–enabled three-dimensional (3D) wireless networks, where the IRS is mounted on aerial platforms, e.g., unmanned aerial vehicles (UAVs), have received significant attention for intelligent signal reflection [22, 19, 14, 41]. Different from traditional terrestrial IRSs, which are inherently restricted to serving only the nodes located within the same half-space of the reflective surface, an aerial IRS can achieve near-omnidirectional reflection and facilitate line-of-sight (LoS) communication between transmitters and receivers due to its elevated position relative to the ground. In [22], transmit beamforming, IRS placement, and 3D passive beamforming were jointly optimized to maximize the worst-case signal-to-noise ratio (SNR) within a target area, utilizing a beam broadening and flattening technique to achieve significant performance gains over the fixed IRS. Furthermore, the maneuverability of UAVs allows the location and orientation of an aerial IRS to be dynamically adjusted, which introduces additional degrees of freedom for system optimization and enables highly flexible 3D network design that can adapt to real-time changes in communication demands. The work [19] developed an efficient alternating optimization algorithm in a UAV-mounted IRS-aided multicast system with the joint design of the IRS’s 3D location, 3D orientation, and passive beamforming. Despite many advantages mentioned above, aerial IRS poses fundamental limitations for several reasons. First, the endurance, stability, and controllability of the aerial platform remain critical technical concerns, as the platform must support sustained operation under varying environmental conditions [26]. Second, platform drift and vibration, which influence reflection accuracy and system stability, as well as the complexity associated with optimizing the placement and trajectory of aerial platforms, impose significant design burdens [7]. Third, regulatory constraints related to flight zones and airspace management may further limit the large-scale or long-term deployment of aerial IRSs in practical scenarios [44].
@@
Considering the ease of implementation provided by the fixed IRS and the high flexibility offered by the aerial IRS, the rotatable IRS has been proposed, which focuses on dynamic orientation adjustment rather than full spatial displacement [40, 8, 18, 39, 13, 38]. To be specific, a rotatable IRS can flexibly vary its rotation angles at a fixed position, thereby dynamically reconfiguring its reception and reflection patterns and providing additional spatial degrees of freedom without physical movement. By properly adjusting its rotation angle, a single rotatable IRS can achieve panoramic coverage or full-angle reflection, enabling it to serve any transmitter-receiver pair within its coverage region. Moreover, the practical implementation of a rotatable IRS resembles existing rotatable antennas, where the IRS is mounted on a motor-driven shaft, and its rotation angles is precisely controlled through mechanical adjustment mechanisms, which can also be considered as a special case of the general six-dimensional movable antenna (6DMA) model with both 3D position and 3D orientation control [32, 33, 34, 48, 21, 25]. Such an architecture ensures strong compatibility with existing wireless infrastructures while avoiding the need for continuous positioning, complex mechanical control, and the high power consumption typically associated with mobile platforms. Because of its simple structure, low energy cost, and low latency, a rotatable IRS is appealing for scenarios that require rapid adaptation but where large-scale mechanical movement is impractical. The rotatable IRS has been extensively investigated under various communication systems, including cellular systems [8], mobile edge computing systems [18], and multicast systems [39]. In [8], the received power was maximized when the incident angle was zero. The work [18] proposed a deep learning-based algorithm for the joint design of IRS orientation, phase shifts, and offloading strategies. The base station (BS) beamforming, IRS phase shifts, and orientation adjustment were jointly optimized to maximize the multicast rates in [18].
@@
Despite the effectiveness of the IRS rotation strategies, most existing studies have relied on a simplified and idealized channel model that primarily accounts for the propagation path loss and orientation parameters, while overlooking other crucial physical-layer characteristics such as the effective reception area of the IRS, as well as its reception and reflection efficiency [3, 2, 28]. Consequently, the corresponding deployment strategies may underperform in practical scenarios with non-ideal channels, as they fail to capture the geometric and physical constraints inherent in practical channel conditions [4]. To address this issue, the work [36] assumed an IRS with predetermined electromagnetic properties and focused on how its steering capability affects the spatial power redistribution of the reflected beam. Furthermore, a physics‑based model was proposed in [24], where the IRS elements were grouped into tiles whose effects on the wireless channel were characterized by solving the electric and magnetic field integral equations. In the above studies, the IRS is represented as a continuous reflecting surface whose global response determines the scattered beam. It provides analytical simplicity and facilitates theoretical investigations. In contrast, the IRS can be modeled as a planar distribution of discrete controllable elements that capture the collective response produced by the individual reflecting elements, which provides a more accurate and practical framework for IRS design by capturing its physical characteristics. One of the main advantages lies in allowing the electromagnetic properties to be directly integrated with circuit‑level models of the tunable elements, thereby bridging the gap between idealized analysis and practical design. In [37], the authors evaluated the scattering performance, power, and size of a single IRS element and then maximized the received signal power toward a specific user. Furthermore, the rotation of the IRS was optimized to maximize the worst SNR over all locations in a target area in both indoor [20] and outdoor scenarios [46]. However, in [37, 20, 46], the radiation patterns of the transmit and receive antennas, as well as the IRS elements, are modeled with angle‑dependent factors derived from a simplified cosine law, which may lead to limitations of rotation optimization in terms of SNR.
@@ II System Model
As shown in Fig. 1, we consider an IRS-aided wireless communication system, where the IRS is deployed to assist from the multi-antenna BS to the single-antenna user within the target area $\mathcal{A}$. We focus on a challenging scenario where the direct BS-user link is blocked by obstacles and assume that all the considered channels follow the free-space LoS model [10, 15]. The IRS consists of a total of $N=N_{\mathrm{c}}\times N_{\mathrm{r}}$ elements, arranged in $N_{\mathrm{r}}$ rows and $N_{\mathrm{c}}$ columns, with each element have the size of $\bar{l}\times\bar{l}$. The BS is equipped with a conventional UPA placed on the $x$-$y$ plane, and the total number of antennas is $M=M_{\mathrm{c}}\times M_{\mathrm{r}}$ antennas, where $M_{\mathrm{r}}$ and $M_{\mathrm{c}}$ denote the number of antennas along the $x$- and $y$-axis, respectively. In the considered system, we establish a global Cartesian coordinate system (CCS). Assuming that the target area $\mathcal{A}$ is a rectangular area with length $A_{\mathrm{x}}$ and width $A_{\mathrm{y}}$ on the $x$-$y$ plane, its center is denoted as $\bm{p}_{\mathrm{r}}=\left[x_{\mathrm{r}},y_{\mathrm{r}},0\right]^{T}$. Thus, any user location within the target area $\mathcal{A}$ is located at $\bm{p}_{\mathrm{U}}=\left[x_{\mathrm{U}},y_{\mathrm{U}},0\right]^{T}$ with $x_{\mathrm{U}}\in\left[x_{\mathrm{r}}-A_{\mathrm{x}}/2,x_{\mathrm{r}}+A_{\mathrm{x}}/2\right]$ and $y_{\mathrm{U}}\in\left[y_{\mathrm{r}}-A_{\mathrm{y}}/2,y_{\mathrm{r}}+A_{\mathrm{y}}/2\right]$.
@@
We denote $\varpi_{\mathrm{B}}^{\mathrm{c}}=\arccos\left({\frac{{({{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}})\cdot{\bm{m}}_{\mathrm{B}}^{\mathrm{c}}}}{{\|{{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}}\|\cdot\|{{\bm{m}}_{\mathrm{B}}^{\mathrm{c}}}\|}}}\right)$, $\varpi_{\mathrm{B}}^{\mathrm{r}}=\arccos\left({\frac{{({{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}})\cdot{\bm{m}}_{\mathrm{B}}^{\mathrm{r}}}}{{\|{{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}}\|\cdot\|{{\bm{m}}_{\mathrm{B}}^{\mathrm{r}}}\|}}}\right)$, ${\varpi_{\mathrm{c}}}=\arccos\left({\frac{{({{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}})\cdot{\bm{m}}_{\mathrm{c}}}}{{\|{{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}}\|\cdot\|{{\bm{m}}_{\mathrm{c}}}\|}}}\right)$, and ${\varpi_{\mathrm{r}}}=\arccos\left({\frac{{({{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}})\cdot{\bm{m}}_{\mathrm{r}}}}{{\|{{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}}\|\cdot\|{{\bm{m}}_{\mathrm{r}}}\|}}}\right)$ as the angles between $({{{\bm{p}}_{{\mathrm{B,1}}}}-{{\bm{p}}_{1}}})$ and ${\bm{m}}_{\mathrm{B}}^{\mathrm{c}}$}, ${\bm{m}}_{\mathrm{B}}^{\mathrm{r}}$}, ${\bm{m}}_{\mathrm{c}}$}}, and ${\bm{m}}_{\mathrm{r}}$}}, respectively. According to [9], the LoS channel from the BS to the IRS can be assumed to be a rank-one matrix if $t_{1,1}\geq\sqrt{N}l^{2}/\lambda$. To facilitate the design of the BS and IRS beamforming, we provide a tractable characterization for the BS-IRS channel in the following proposition.
@@
Given $\bm{\Phi}$ and $\bm{\Omega}$, the optimal transmit beamforming at the BS is given by [22]
@@ VI Conclusion
We consider a perfectly conducting rectangular plate with dimensions $l_{1}\times l_{2}$ and negligible thickness, located in the xyx-y plane. A distant point source radiates linearly polarized waves with wave number $k=\frac{2\pi}{\lambda}$. As shown in Fig. 13, let $\theta^{\mathrm{i}}$, $\phi^{\mathrm{i}}$, $\theta^{\mathrm{r}}$, and $\phi^{\mathrm{r}}$ denote the azimuth and elevation angles of the incident wave of the center of plane, as well as the elevation and azimuth angles of the reflected wave of the center of plane. The electric and magnetic field components of the incident wave is given by [23, 27]